OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (–7, 8) and (0, –8).

OpenStudy (anonymous):

first find the slope

OpenStudy (anonymous):

which is -16/7

OpenStudy (anonymous):

good next use point-slope form to find the equations Point-Slope Form: y-y1 = m(x-x1)

OpenStudy (anonymous):

I got y=-16/7x-1

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

would standard form be 16/7x+y=-1